Two parallel wire carry equal current of 10A along same direction and are seperated by distace of 2cm find magnetic field at point which is 2cn away from each wire??
 
 
 
 

Grade:

1 Answers

Arun
5 years ago

Dear studet

 

B = mo * I / 2 * pi * r = 10^-7 * 2 * 10 / 2 * 10^-2

 

B = 10^-4 T

So net is [(10–4)2 + (10–4)2 + 2(10–8) Cos 60°]1/2

= 10–4[1 + 1 + 2 × ½ ]1/2 = 10-4 × 3 T = 1.732 × 10–4T